# On Nontrivial Winning and Losing Parameters of Schmidt Games

### v1

@@ 3. General strategies in the Schmidt’s game
If $A(I)$ was Alice’s winning positional strategy on the set $S$, then $A(I-\xi)+\xi$ is a winning positional strategy on the set $D$. ∎
Now, we use the lemmas we have to prove the following statement:
Theorem 3.10.
Fix $m\in\mathbb{Z}$. Let $D\subseteq\mathbb{R}$ be some set with the following properties:
  • $m^{k}D=m^{k}+D=D$ for any integer $k$.
  • $D\sqcup(-D)\sqcup Q=\mathbb{R}$, where $Q\subseteq\mathbb{Q}$ is some set.
  • $D+\xi\subseteq D$ for some $\xi\notin\mathbb{Q}$.
Then:
  1. (1)
    If $\alpha\geq\beta$, then the set $D$ is $(\alpha,\beta)$-losing.
  2. (2)
    If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then the set $D$ is $(\alpha,\beta)$-winning.
Proof.
  1. (1)
    follows from Lemma 3.8.
  2. (2)
    By Lemma 3.8 we know that the set $D\sqcup Q$ is $(\alpha,\beta)$-winning. As $Q\subseteq\mathbb{Q}$, one has $D\sqcup Q+\xi\subseteq D$ for any $\xi\notin\mathbb{Q}$. By Lemma 3.9 then, $D$ is $(\alpha,\beta)$-winning for $\beta>\alpha$.
Corollary 3.11.
Let $D\subseteq\mathbb{R}$ be some set with the following properties:
  • $m_{i}^{k}D=m_{i}^{k}+D=D$ for any integer $k$ and some $m_{1},\,m_{2}\in\mathbb{N}$ with property $\log_{m_{1}}(m_{2})\notin\mathbb{Q}$;
  • $D\sqcup(-D)\sqcup Q=\mathbb{R}$, where $Q\subseteq\mathbb{Q}$ is some set;
  • $D+\xi\subseteq D$ for some $\xi\notin\mathbb{Q}$.
If $\beta>\alpha$, then the set $D$ is $(\alpha,\beta)$-winning.
Indeed: for any parameters $(\alpha,\beta)$ in this case either $\log_{m_{1}}(\alpha\beta)$ or $\log_{m_{2}}(\alpha\beta)$ is irrational, so by Theorem 3.10 $D$ is $(\alpha,\beta)$-winning.
4. Example: sets with a nontrivial diagram
Theorem 3.10 and Corollary 3.11 give the complete description of the nontrivial Schmidt diagram for sets with special properties. The goal of this section is to give an example of a set satisfying these properties, thus proving Theorem 2.1.
Fix $\xi\notin\mathbb{Q}$. Then, we define an auxiliary set $\Xi$ by $\Xi=\{r+q\xi:\,\,r,q\in\mathbb{Q},\,q>0\}$. One can see that $-\Xi=\{r+q\xi:\,\,r,q\in\mathbb{Q},\,q<0\}$ is disjoint with $\Xi$, and one has $\Xi+\xi\subseteq\Xi$ by construction.
Now, let us focus on the set $E=\mathbb{R}\setminus\Big{(}\mathbb{Q}\cup\Xi\cup(-\Xi)\Big{)}$. We will divide it into the union of the sets $\overline{x}=\{r+q\xi+px:\,\,r,q,p\in\mathbb{Q},\,\,p\neq 0\}$ over $x\in E$. Easy to see that $\overline{x}$ is the equivalence class under relation $\sim$ defined as follows: $x\sim y$ if $y=r+q\xi+px$ with the restrictions as above, therefore $E$ is the disjoint union of the sets $\overline{x}$ over pairwise non-equivalent $x$.
Next, let $\underline{x}=\{r+q\xi+px:\,\,r,q,p\in\mathbb{Q},\,\,p>0\}$. Then $\overline{x}=\underline{x}\sqcup\underline{-x}$. Indeed: if $\underline{x}\,\text{and}\,\underline{-x}$ intersect, then $x\in\Xi\sqcup(-\Xi)$, which we do not consider. Finally, note that $\underline{-x}=-\underline{x}$.
Let us enumerate all the sets $\overline{x}$ by some index set $\mathscr{A}$ and denote the corresponding set $E_{\alpha}$ where $\alpha\in\mathscr{A}$. Choose some $x_{\alpha}\in A_{\alpha}$ and denote $\underline{x_{\alpha}}=C_{\alpha}$; then $C_{\alpha}\sqcup(-C_{\alpha})=E_{\alpha}$. Therefore for $C=\bigsqcup\limits_{\alpha\in\mathscr{A}}C_{\alpha}$ one has $C\sqcup(-C)=E$; by construction also $C+\xi_{j}=C$ (as this works for each of $\underline{x}$).
Finally, we define $\mathscr{S}=\mathscr{S}(\xi)=C\sqcup\Xi$. Then $-\mathscr{S}=(-C)\sqcup(-\Xi)$, and one has $\mathscr{S}\sqcup(-\mathscr{S})\sqcup\mathbb{Q}=\mathbb{R}$. One also has:
  • $m^{k}\mathscr{S}=m^{k}+\mathscr{S}=\mathscr{S}$ for any integer $k$ and any $m\in\mathbb{N}$;
  • $\mathscr{S}+\xi\subseteq\mathscr{S}$.
Therefore, by Theorem 3.10 and Corollary 3.11, $\mathscr{S}$ is $(\alpha,\beta)$-winning for $\beta>\alpha$ and $(\alpha,\beta)$-losing for $\beta\leq\alpha$. This proves Theorem 2.1.
5. Digit frequencies
Next, we will investigate a separate but similar construction regarding frequencies of digits in base-$2$ expansions of real numbers. It is known that these sets are closely related to density and dimension, and hence they have been studied in the past (see [V59, FFNS21]).
First, some setup. Consider the base-$2$ expansions of the form $x=x_{0}.x_{1}x_{2}\cdots$ where $x_{0}$ is an integer and $x_{i}\in\{0,1\}$ are the digits in the base-$2$ expansion of $x$. In this paper, if $x$ can be represented by two seperate expansions, we take the lexicographically minimal expansion (for example, $0.100\ldots$ would instead be written as $0.011\ldots$). We define
and similarly,
Define the sets
$D^{+}_{c}$ :=$:=\{x\in{\mathbb{R}}:d^{+}(x,0)>c\},$
$D^{-}_{c}$ :=$:=\{x\in{\mathbb{R}}:d^{-}(x,0)>c\}.$
The following is known about the Schmidt diagrams of these sets.
Theorem 5.1.
[FFNS21] The set $D^{+}_{c}$ is winning for $\log(8\alpha)/\log(\alpha\beta)>c$ and losing for $\log(\alpha/8)/\log(\alpha\beta)\leq c$.
The results of our research expand on that bound. To do so, we require the following lemma, known as the Equidistribution Theorem:
@@ 5. Digit frequencies
Theorem 5.3.
Suppose $\log_{2}{\alpha\beta}\notin\mathbb{Q}$. Then, the sets $D^{\pm}_{c}$ are winning for $\alpha\leq\frac{1}{4}$ and $\log(4\alpha)/\log(\alpha\beta)>c$, and losing for $\beta\leq\frac{1}{4}$ and $\log(\alpha/4)/\log(\alpha\beta)\leq c$.
In the special case the bounds given by Theorem 5.3 are presented on Figure 5.
@@
We will show the winning strategy for Alice.
Suppose Bob begins by playing an interval with radius $r_{0}$. Alice will play in the following way: suppose Bob plays some interval $B_{t}$. Define $k_{t}=-\left\lfloor\log_{2}{(\alpha\beta)^{t}r_{0}}\right\rfloor$, and pick the smallest $m\in\mathbb{Z}$ such that $m/2^{k_{t}}\in B_{t}$. Then, it follows that the dyadic interval $\left(\frac{m}{2^{k_{t}}},\frac{m+1}{2^{k_{t}}}\right)\subset B_{t}$. Then, Alice plays $A_{t}=\left(\frac{m}{2^{k_{t}}},\frac{m}{2^{k_{t}}}+2\alpha(\alpha\beta)^{t}r_{0}\right)$. Define $l_{t}$ such that $k_{t}+l_{t}=-\left\lfloor\log_{2}{\alpha(\alpha\beta)^{t}r_{0}}\right\rfloor-2$. Then, Alice ensures that for any $x\in A_{t}$, the $(k_{t}+1)$th digit to the $(k_{t}+l_{t})$th digit is equal to 0. In other words, she guarantees $l_{t}$ zeroes on this turn. For convenience, let $a=\log_{2}\alpha$, $b=\log_{2}\beta$, and $\rho=\log_{2}r_{0}$; note that $a+b\notin\mathbb{Q}$. We have
@@
This implies that by the Equidistribution Theorem,
@@
Now, notice that for $k_{n}\leq k<k_{n+1}$ one has
@@
$=\frac{\log(4\alpha)}{\log(\alpha\beta)}>c$
as desired. Therefore, Alice wins. The losing condition is analogous, so we omit it.
For the proofs below, we will use the following
@@
Let $S\subseteq\mathbb{N}$ be an infinite set of density zero. Let $d^{\pm}(x,0)=c$; then
$d^{\pm}\left(x+\sum\limits_{j\in S}\frac{1}{2^{j}},0\right)\geq c.$
@@
For $x\in\mathbb{Z}[\frac{1}{2}]$ one has $d^{\pm}t(x+\sum\limits_{j\in S}\frac{1}{2^{j}},0)=d^{\pm}(\sum\limits_{j\in S}\frac{1}{2^{j}},0)=1$, as addition of $x$ changes only finitely many digits. Now assume $x\notin\mathbb{Z}[\frac{1}{2}]$.
@@
$d^{\pm}(x+\sum\limits_{j\in S_{2}}\frac{1}{2^{j}},0)={\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\sqcup S_{2}\}}{k}=$
${\lim\limits_{k\rightarrow\infty}}^{\pm}\Big{(}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-\frac{\#\{1\leq i\leq k:\,i\in S_{2}\}}{k}\Big{)}=$
${\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-\lim\limits_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:\,i\in S_{2}\}}{k}=$
@@
Now it remains to prove the statement only for $S$ such that $S\subseteq Q$. Let us define $R\subseteq\mathbb{N}$ by $\sum\limits_{j\in Q}\frac{1}{2^{j}}+\sum\limits_{j\in S}\frac{1}{2^{j}}=\sum\limits_{j\in R}\frac{1}{2^{j}}$. It is easy to check that:
@@
This immediately implies that $d^{\pm}(x+\sum\limits_{j\in S}\frac{1}{2^{j}},0)\geq d^{\pm}(x,0)=c$ and completes the proof. ∎
@@
We now focus on the special case $c=\frac{1}{2}$. We will use the construction similar to the set $\mathscr{D}$ but connected to digit frequencies.
Let $\mathscr{F}=D^{-}_{\frac{1}{2}}\cup(\mathscr{D}(\xi)\cap\{x:\,d^{+}(x,0)\geq\frac{1}{2}\})$ where $\xi=\sum\limits_{i=1}^{\infty}\frac{1}{2^{2^{i}}}$. We want to show that the set $\mathscr{F}$ satisfies the conditions of Theorem 3.10 for $m=2$.
Proposition 5.5.
@@
    $2^{k}\mathscr{F}=2^{k}+\mathscr{F}=\mathscr{F}$ for any integer $k$;
@@
    $\mathscr{F}\sqcup(-\mathscr{F})\sqcup Q=\mathbb{R}$ where $Q\subseteq\mathbb{Q}$;
  3. (3)
    $\mathscr{F}+\xi\subseteq\mathscr{F}$.
Proof.
  1. (1)
    Follows from the similar properties of all the sets in construction of $\mathscr{F}$;
  2. (2)
    It’s easy to notice that
    =$-\mathscr{F}=\left\{x\notin\mathbb{Z}\left[\frac{1}{2}\right]:\,d^{+}(x,0)<\frac{1}{2}\right\}\cup\left((-\mathscr{D}(\xi))\cap\left\{x:\,d^{-}(x,0)\leq\frac{1}{2}\right\}\right),$
    which shows that $\mathscr{F}$ and $-\mathscr{F}$ are disjoint. If $x\notin\mathscr{F}\sqcup(-\mathscr{F})$, then either $x\in\mathbb{Z}[\frac{1}{2}]$ or $d^{-}(x,0)\leq\frac{1}{2}\leq d^{+}(x,0)$ and $x\notin\mathscr{D}\sqcup(-\mathscr{D})$, which implies $x\in\mathbb{Q}$.
  3. (3)
    By Lemma 5.4 all the sets in the construction of $\mathscr{F}$ satisfy this condition.
This immediately proves that
  • If $\alpha\geq\beta$, then the set $\mathscr{F}$ is $(\alpha,\beta)$-losing;
  • If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then the set $\mathscr{F}$ is $(\alpha,\beta)$-winning.
Let’s also note that $\{x:\,d^{-}(x,0)>\frac{1}{2}\}\subseteq\mathscr{F}\subseteq\{x:\,d^{+}(x,0)\geq\frac{1}{2}\}$, thus
Corollary 5.6.
  1. (1)
    If $\alpha\geq\beta$, then the set $D^{-}_{\frac{1}{2}}=\{x:\,d^{-}(x,0)>\frac{1}{2}\}$ is $(\alpha,\beta)$-losing;
  2. (2)
Conjecture 5.7.
We believe that $D(D^{-}_{c})=\{(\alpha,\beta):\,\log(\alpha)/\log(\alpha\beta)>c\}$.
6. B-Badly Approximable Numbers
It is known that the badly approximable numbers are dense and have measure zero. Interestingly enough, despite having measure zero, Schmidt showed that they are winning for the largest possible nontrivial bounds:
Theorem 6.1.
[S66] The set $\operatorname{BA}$ is winning for all $(\alpha,\beta)\in\hat{D}$; that is, $D(\operatorname{BA})=\hat{D}$.
In a similar vein to the badly approximable numbers investigated by Schmidt, we define the set of B-Badly Approximable Numbers as
$\operatorname{B-BA}:=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{B^{k}}\right|>\frac{c}{B^{k}}\text{ for some $c>0$ and all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}$
and a closely related set
$\operatorname{B-BA}(c):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{B^{k}}\right|>\frac{c}{B^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}.$
These definitions imply $\operatorname{B-BA}=\bigcup_{c>0}\operatorname{B-BA}(c)$. We note that the $\operatorname{B-BA}$ set has been previously studied in [N10] and [PS08].
The following statement follows from work of Schmidt:
Theorem 6.2.
[S66] The set $\operatorname{2-BA}$ is winning for all $(\alpha,\beta)\in\hat{D}$; that is, $D(\operatorname{2-BA})=\hat{D}$.
While both $\operatorname{2-BA}$ and $\operatorname{2-BA}(c)$ did not produce any new Schmidt diagrams, there is a simple change we can make that will produce both nontrivial winning and nontrivial losing zones. For any integer $B\geq 2$, define
$\operatorname{\widetilde{B-BA}}(c,N):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{B^{k}}\right|>\frac{c}{B^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\ \text{s.t.}\ k>N\right\}$
and
$\operatorname{\widetilde{B-BA}}(c):=\bigcup_{N\in{\mathbb{N}}}\operatorname{\widetilde{B-BA}}(c,N).$
Throughout this chapter, we define $\gamma_{A}=1-2\alpha+\alpha\beta$ and $\gamma_{B}=1-2\beta+\alpha\beta$. We begin results in this section with a useful general lemma:
@@ 6. B-Badly Approximable Numbers
Suppose Bob plays any interval $B_{t}$ and $\gamma_{A}>0$. Denote the center of $B_{t}$ by $c_{t}$. Then, Bob can always choose an interval $B_{t+1}$ such that $|c_{t+1}-c_{t}|\leq\frac{|B_{t}|}{2}\gamma_{A}$. As a corollary, Bob can guarantee that for any $m\in\mathbb{N}$, $|c_{t+m}-c_{t}|\leq\frac{|B_{t}|}{2(1-\alpha\beta)}\gamma_{A}$. Analogous results hold for Alice, with $\gamma_{A}$ replaced by $\gamma_{B}$.
@@
Let’s denote center of $B_{t}$ by $c_{t}$. It is clear that Bob can always choose such a $B_{t+1}$ that $|c_{r+1}-c_{r}|\leq\frac{|B_{t}|}{2}\gamma_{A}$, since no matter which side Alice picks her interval on, Bob will go the opposite way. Extending this logic, it follows that $|c_{m+r}-(a_{0}+k\delta)|\leq\frac{\rho(\alpha\beta)^{m}}{2}\gamma_{A}(1+(\alpha\beta)+(\alpha\beta)^{2}+\ldots)=\frac{\rho(\alpha\beta)^{m}}{2(1-\alpha\beta)}\gamma_{A}$, which completes the proof.
The following theorems provide the winning and losing conditions for $\operatorname{\widetilde{B-BA}}$ sets.
Theorem 6.4.
If
$\frac{2c(1-\beta)(1-\alpha\beta)}{\beta\gamma_{A}}>1,$
then for any $B$, the set $\operatorname{\widetilde{B-BA}}(c)$ is $(\alpha,\beta)$-losing.
Proof.
@@
$\operatorname{\widetilde{B-BA}}(c)^{c}=\left\{x:\,\,\forall N\,\,\exists\,n>N\,\,\text{and}\,\,k\,\in\,\mathbb{Z}:\,\,\left|x-\frac{k}{B^{n}}\right|<\frac{c}{B^{n}}\right\}=$
$=\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{n=N}^{\infty}\left(\bigcup\limits_{k\in\mathbb{Z}}\left[\frac{k}{B^{n}}-\frac{c}{B^{n}},\frac{k}{B^{n}}+\frac{c}{B^{n}}\right]\right)\subseteq\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{n=N}^{\infty}\left(\bigcup\limits_{k\in\mathbb{Z}}\left(\frac{k}{B^{n}}-\frac{c}{B^{n}},\frac{k}{B^{n}}+\frac{c}{B^{n}}\right)\right),$
and let $J_{n}=\bigcup\limits_{k\in\mathbb{Z}}\left(\frac{k}{B^{n}}-\frac{c}{B^{n}},\frac{k}{B^{n}}+\frac{c}{B^{n}}\right)$. Fix $\varepsilon>0$ such that
$\varepsilon=\frac{2c(1-\beta)(1-\alpha\beta)}{\beta\gamma_{A}}-1.$
@@
    $\frac{1}{B}\left(\frac{2c(1-\alpha\beta)}{\gamma_{A}}\right)$
    centered at $1/B$. By Lemma 1.2, Bob can play in such a way that in finitely many moves, $B_{s}\subset J_{1}$.
@@
    $\log_{B}\left(\frac{1}{1-\beta}\right)$$\leq\left\{\left(t\cdot\log_{B}\alpha\beta\right)+\log_{B}|A_{0}|\right\}$ (6.1)
    $\leq\log_{B}\left(\frac{1}{1-\beta}\right)+\log_{B}(1+\varepsilon),$

### v3

@@ 3. General strategies in the Schmidt game
If $A(I)$ was Alice’s winning positional strategy on the set $E$, then $A(I-\xi)+\xi$ is a winning positional strategy on the set $S$. ∎
4. Proof of Theorem 2.2
We start with the following classical statement about uniform distribution:
@@
We will prove two separate statements, which together imply the result of Theorem 2.2: Proposition 4.2 and Proposition 4.3.
Proposition 4.2.
Let $c\in(0,1)$, and suppose $\log_{2}{\alpha\beta}\notin\mathbb{Q}$. Then, the sets $F^{\pm}_{c}$ are $(\alpha,\beta)$-winning if $\log(4\alpha)/\log(\alpha\beta)>c$, and $(\alpha,\beta)$-losing if $\log(\alpha/4)/\log(\alpha\beta)\leq c$.
@@
First, let us notice that our winning conditions imply that $\alpha<\frac{1}{4}$, and the losing conditions imply that $\beta\leq\frac{1}{4}$. We will show the winning strategy under these conditions for Alice in the $(\alpha,\beta)$-game on the set $F^{-}_{c}$ (and thus on $F^{+}_{c}$, since $F^{-}_{c}\subseteq F^{+}_{c}$).
Let $r_{0}$ be the radius of $B_{0}:\,\,r_{0}=\frac{1}{2}|B_{0}|$, and suppose Bob played some interval $B_{t}$. Define $k_{t}=-\left\lfloor\log_{2}{(\alpha\beta)^{t}r_{0}}\right\rfloor$. Consider the smallest $m\in\mathbb{Z}$ such that $m/2^{k_{t}}\in B_{t}$. For such an $m$, one has $\left[\frac{m}{2^{k_{t}}},\frac{m+1}{2^{k_{t}}}\right]\subset B_{t}$. Then, Alice plays $A_{t}=\left[\frac{m}{2^{k_{t}}},\frac{m}{2^{k_{t}}}+2\alpha(\alpha\beta)^{t}r_{0}\right]$. Define $l_{t}$ such that $k_{t}+l_{t}=-\left\lfloor\log_{2}{\alpha(\alpha\beta)^{t}r_{0}}\right\rfloor-2$. Then, Alice ensured by her move that for any $x\in A_{t}$, the $(k_{t}+1)$-th to the $(k_{t}+l_{t})$-th digits are equal to 0. In other words, she guarantees $l_{t}$ zeroes on this turn. For convenience, let $a=\log_{2}\alpha$, $b=\log_{2}\beta$, and $\rho=\log_{2}r_{0}$; note that $a+b\notin\mathbb{Q}$, and $a+2<0$. We have
@@
This implies that by the Lemma 4.1,
@@
Now, let $x=x_{0}+\sum\limits_{i=1}^{\infty}x_{i}2^{-i}=\bigcap\limits_{t=0}^{\infty}A_{t}$ be the outcome of the game. We notice that for $k_{n}\leq k<k_{n+1}$ one has
@@
$=\frac{\log(4\alpha)}{\log(\alpha\beta)}>c,$
which means that $x\in F^{-}_{c}$. Therefore, Alice wins. The losing condition is analogous, so we omit it.
For the next statement, we fix $c=\frac{1}{2}$. Together with the proposition above, it clearly implies the statement of Theorem 2.2 due to trivial inclusions: $F_{c_{1}}^{\pm}\subseteq F_{c_{2}}^{\pm}$ if $c_{1}\geq c_{2}$.
Proposition 4.3.
  1. (1)
    If $\alpha\geq\beta$, then the set $F^{-}_{\frac{1}{2}}=\{x:\,d^{-}(x,0)>\frac{1}{2}\}$ is $(\alpha,\beta)$-losing;
  2. (2)
In order to show Proposition 4.3, we will need the following observation:
@@
Let $\mathfrak{S}\subseteq\mathbb{N}$ be an infinite set of density zero, and define $\xi:=\sum\limits_{j\in\mathfrak{S}}\frac{1}{2^{j}}$. Suppose $d^{\pm}(x,0)=c$; then
$d^{\pm}\left(x+\xi,0\right)\geq c.$
@@
For $x\in\mathbb{Z}[\frac{1}{2}]$ one has $d^{\pm}t(x+\sum\limits_{j\in\mathfrak{S}}\frac{1}{2^{j}},0)=d^{\pm}(\sum\limits_{j\in\mathfrak{S}}\frac{1}{2^{j}},0)=1$, as addition of $x$ changes only finitely many digits. Now assume $x\notin\mathbb{Z}[\frac{1}{2}]$.
@@
$d^{\pm}(x+\sum\limits_{j\in\mathfrak{S}_{2}}\frac{1}{2^{j}},0)={\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\sqcup\mathfrak{S}_{2}\}}{k}=$
${\lim\limits_{k\rightarrow\infty}}^{\pm}\Big{(}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-\frac{\#\{1\leq i\leq k:\,i\in\mathfrak{S}_{2}\}}{k}\Big{)}=$
${\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-\lim\limits_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:\,i\in\mathfrak{S}_{2}\}}{k}=$
@@
Now it remains to prove the statement only for $\mathfrak{S}$ such that $\mathfrak{S}\subseteq Q$. Let us define $R\subseteq\mathbb{N}$ by $\sum\limits_{j\in Q}\frac{1}{2^{j}}+\sum\limits_{j\in\mathfrak{S}}\frac{1}{2^{j}}=\sum\limits_{j\in R}\frac{1}{2^{j}}$. It is easy to check that:
@@
This immediately implies that $d^{\pm}(x+\sum\limits_{j\in\mathfrak{S}}\frac{1}{2^{j}},0)\geq d^{\pm}(x,0)=c$ and completes the proof. ∎
@@
We are ready to prove Proposition 4.3.
Proof of Proposition 4.3. Let $S=F_{\frac{1}{2}^{-}}\setminus\mathbb{Z}[\frac{1}{2}]$. Then, using the properties 2 and 3 listed above, it is easy to verify that $-S=\{x:\,\,d^{+}(-x,0)<\frac{1}{2}\}\setminus\mathbb{Z}[\frac{1}{2}]$ is a set disjoint with $S$. Let us define $Q$ under the condition $S\sqcup(-S)\sqcup Q=\mathbb{R}$; in this case, $S\sqcup Q=\{x:\,d^{+}(x,0)\geq\frac{1}{2}\}\cup\mathbb{Z}[\frac{1}{2}]$.
Property 1 shows that the sets $S$ and $Q$ satisfy the conditions of the Lemma 3.8 with $m=2$. As in addition our sets are Borel, the $(\alpha,\beta)$-game on them is determined, and we can claim that
  • If $\alpha\geq\beta$, then $(\alpha,\beta)$ is a losing point for $S$;
  • If $\beta>\alpha$ and $\log_{2}(\alpha\beta)\notin\mathbb{Q}$, then $(\alpha,\beta)$ is a winning point for $S\sqcup Q$.
Now let us fix $\xi$ as in Lemma 4.4: $\xi$ is irrational, so by the aforementioned lemma $F_{\frac{1}{2}}^{-}+\xi\subseteq S$ and $S\sqcup Q+\xi\subseteq\{x:\,d^{+}(x,0)\geq\frac{1}{2}\}$. The following two observations complete the proof of Proposition 4.3:
@@
    Suppose $F_{\frac{1}{2}}^{-}$ is $(\alpha,\beta)$-winning for $\alpha\geq\beta$; then, by Lemma 3.10, $S$ should also be $(\alpha,\beta)$-winning, which is not true. Provided that the game is determined on $F_{\frac{1}{2}}^{-}$, we conclude that the set $F_{\frac{1}{2}}^{-}$ is $(\alpha,\beta)$-losing.
@@
    Follows directly from Lemma 3.10 and observations above.
5. Proofs of the statements regarding d-Badly Approximable Numbers
We begin results in this section with a useful general lemma:
@@ 5. Proofs of the statements regarding d-Badly Approximable Numbers
Suppose Bob plays any interval $B_{t}$ and $\gamma>0$. Denote the center of $B_{t}$ by $c_{t}$. Then, Bob can always choose an interval $B_{t+1}$ such that $|c_{t+1}-c_{t}|\leq\frac{|B_{t}|}{2}\gamma$. As a corollary, Bob can guarantee that for any $m\in\mathbb{N}$, $|c_{t+m}-c_{t}|\leq\frac{|B_{t}|}{2(1-\alpha\beta)}\gamma$. Analogous results hold for Alice, with $\gamma$ replaced by $1-2\beta+\alpha\beta$.
@@
Let’s denote center of $B_{t}$ by $c_{t}$. It is clear that Bob can always choose such a $B_{t+1}$ that $|c_{r+1}-c_{r}|\leq\frac{|B_{t}|}{2}\gamma$, since no matter which side Alice picks her interval on, Bob will go the opposite way. Extending this logic, it follows that $|c_{m+r}-(a_{0}+k\delta)|\leq\frac{\rho(\alpha\beta)^{m}}{2}\gamma(1+(\alpha\beta)+(\alpha\beta)^{2}+\ldots)=\frac{\rho(\alpha\beta)^{m}}{2(1-\alpha\beta)}\gamma$, which completes the proof.
Now we are ready to give the proofs of our main theorems.
Proof of Theorem 2.5
@@
$\operatorname{\widetilde{d-BA}}(c)^{c}=\left\{x:\,\,\forall N\,\,\exists\,n>N\,\,\text{and}\,\,k\,\in\,\mathbb{Z}:\,\,\left|x-\frac{k}{d^{n}}\right|\leq\frac{c}{d^{n}}\right\}=$
$=\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{n=N}^{\infty}\left(\bigcup\limits_{k\in\mathbb{Z}}\left[\frac{k}{d^{n}}-\frac{c}{d^{n}},\frac{k}{d^{n}}+\frac{c}{d^{n}}\right]\right)\subseteq\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{n=N}^{\infty}\left(\bigcup\limits_{k\in\mathbb{Z}}\left(\frac{k}{d^{n}}-\frac{c}{d^{n}},\frac{k}{d^{n}}+\frac{c}{d^{n}}\right)\right),$
and let $J_{n}=\bigcup\limits_{k\in\mathbb{Z}}\left(\frac{k}{d^{n}}-\frac{c}{d^{n}},\frac{k}{d^{n}}+\frac{c}{d^{n}}\right)$. Fix $\varepsilon>0$ such that
$\varepsilon=\frac{2c(1-\beta)(1-\alpha\beta)}{\beta\gamma}-1.$
@@
    $\frac{1}{d}\left(\frac{2c(1-\alpha\beta)}{\gamma}\right)$
    centered at $1/d$. By Lemma 1.2, Bob can play in such a way that in finitely many moves, $B_{s}\subset J_{1}$.
@@
    $\log_{d}\left(\frac{1}{1-\beta}\right)$$\leq\left\{\left(t\cdot\log_{d}\alpha\beta\right)+\log_{d}|A_{0}|\right\}$ (5.1)
    $\leq\log_{d}\left(\frac{1}{1-\beta}\right)+\log_{d}(1+\varepsilon),$